Joe has a collection of nickles and dimes worth $6.20. If the number of dimes was doubled and the number of nickles was decreased by 37, the value of the coins would be $8.85. How many nickles and dimes does he have?
Answer by ewatrrr(10682)

 
Hi,
Let x and y represent the number of nickels and dimes respectively
Question states***CENTS makes sense
5x + 10y = 620
5(x-7) + 10(2y) = 885
5x + 10y = 620
5x + 10(2y) = 920
-5x - 10y = -620
5x + 10(2y) = 920
10y = 300
y = 30, the number of dimes. Number of nickels is 64.
CHECKING our Answer***
5(x-7) + 10(2y) = 885
5*57 + 10*60 = 285 + 600 = 885 OR $8.85